# On Supports for Graphs of Bounded Genus [^1] Parts of this work appeared in “On Hypergraph Supports” available at https://arxiv.org/abs/2303.16515, which has now been split into two papers. This is the first paper in the series, and the second paper is available at https://arxiv.org/abs/2504.05039v2

### v2

@@ 6.2 Dual Support
In this section, we show that for a cross-free system on a graph of genus $g$, there is a dual support of genus at most $g$. Unlike the primal case however, we are unable to prove that the algorithm implied in the proof of Theorem 8 runs in polynomial time, even if a cross-free embedding of the system is given. The main difficulty is that in the process of vertex bypassing at a vertex $v$, we add additional vertices while subdividing the edges incident to $v$. In the primal case, the newly added red vertices are not maximal vertices, and therefore we do not apply vertex bypassing to them. In the dual case however, we may require repeatedly applying vertex bypassing to newly created vertices. As a consequence, the size of the graph increases and we are not able to bound the size of the graph by a polynomial.
We show that for constructing a dual support, we can assume that there are no two $H,H^{\prime}\in\mathcal{H}$ such that $V(H)\subseteq V(H^{\prime})$.
@@
Let $(G,\mathcal{H})$ be a graph system. Let $\mathcal{H}^{\prime}\subseteq\mathcal{H}$ be maximal such that $\forall\;H,H^{\prime}\in\mathcal{H}^{\prime}$, $H\setminus H^{\prime}\neq\emptyset$, and that for each $H\in\mathcal{H}\setminus\mathcal{H}^{\prime}$, $H\subseteq H^{\prime}$ for some $H^{\prime}\in\mathcal{H}^{\prime}$. If $Q^{\prime}$ is a dual support for $(G,\mathcal{H}^{\prime})$ then, there is a dual support $Q^{*}$ for $(G,\mathcal{H})$ such that if $Q^{\prime}$ has genus $g$, then $Q^{*}$ has genus $g$.
@@
Consider the containment order $\mathcal{P}=(\mathcal{H},\preceq)$ where for $H,H^{\prime}\in\mathcal{H}$, $H\preceq H^{\prime}$ iff $V(H)\subseteq V(H^{\prime})$. Then $\mathcal{P}$ is a partial ordered set. Let $\ell(\mathcal{P})$ be the maximum length of a chain in $\mathcal{P}$, and $c(\mathcal{P})$ be the number of chains of length $\ell(\mathcal{P})$. We prove by induction on $(\ell(\mathcal{P}),c(\mathcal{P}))$, that a dual support $Q^{\prime}$ for a lexicographically smaller system can be extended to a dual support $Q^{*}$ of a larger system so that $Q^{*}$ has the same genus as $Q^{\prime}$. If we show this, then we can remove all containments, construct a support, and extend it to obtain a dual support for the original system without increasing the genus.
@@
Consider a longest chain $C$ in $\mathcal{P}$, and let $H$ be the minimum element in $C$. Let =$\mathcal{H}^{\prime}=\mathcal{H}\setminus\{H\}$. Then, for containment order $\mathcal{P}^{\prime}$ on $\mathcal{H}^{\prime}$, $(\ell(\mathcal{P}^{\prime}),c(\mathcal{P}^{\prime}))$ is lexicographically smaller than $(\ell(\mathcal{P}),c(\mathcal{P}))$. Therefore, by the inductive hypothesis, the statement of the proposition holds for $(G,\mathcal{H}^{\prime})$, and let $Q^{\prime}$ denote its dual support. To obtain a dual support $Q^{*}$ for $(G,\mathcal{H})$, we add a vertex corresponding to $H$. Let $H^{\prime}$ be an immediate successor of $H$ in $\mathcal{P}$ arbitrarily chosen. Connect $H$ to $H^{\prime}$. Since we added a new vertex of degree 1, it follows that if $Q^{\prime}$ has genus $g$, then $Q^{*}$ has genus $g$.
@@
For each vertex $v\in V(G)$, the subgraphs containing $v$ trivially induce a connected subgraph in $Q^{*}$ since $10pt(v)=$ 1 for all $v\in V(G)$. Consider a special edge $e=\{u,v\}\in E(G)$. Since both end-points belong to a subgraph in $\mathcal{H}$, and $\mathcal{H}_{u}\cap\mathcal{H}_{v}=\emptyset$, it implies that $\mathcal{H}_{u}\not\subseteq\mathcal{H}_{v}$ and $\mathcal{H}_{v}\not\subseteq\mathcal{H}_{u}$. In the construction of $Q^{*}$ therefore, $e$ is not contracted. Further, $u$ and $v$ are contracted to vertices corresponding to the unique subgraph containing $u$ and $v$, respectively. Therefore, $Q^{*}$ satisfies the special edge property. ∎
@@
Therefore, we can assume that for a maximum depth vertex $v$ and each $e\sim v$, $\mathcal{H}_{e}\subset\mathcal{H}_{v}$. We apply VB$(v)$ to obtain the system $(G^{\prime},\mathcal{H}^{\prime})$. By Lemma  25, the new vertices $u_{0},\ldots,u_{deg(v)-1}$ in $G^{\prime}$ obtained on applying Vertex Bypassing to $v$ have depth at most $d-1$. Hence, $(d^{\prime},n_{d^{\prime}})\prec(d,n_{d})$, where $d^{\prime}$ and $n_{d^{\prime}}$ are respectively, the depth of a maximum depth vertex, and the number of vertices of maximum depth in $(G^{\prime},\mathcal{H}^{\prime})$. We can assume that $\mathcal{H}^{\prime}_{u_{i}}\neq\emptyset$ for each $i\in\{0,\ldots,\deg(v)-1\}$. Otherwise, we can contract the edge $\{u_{i},u_{i+1}\}$ (with indices taken $\mathrm{mod}$ $\deg(v)$) as this does not violate the cross-free condition on $\mathcal{H}^{\prime}$. Further, there is an injective correspondence between the special edges in $G$ and the special edges in $G^{\prime}$: For $i=0,\ldots,deg(v)-1$, if $\{v,v_{i}\}$ in $G$ is a special edge, the corresponding edge $\{u_{i},v_{i}\}$ is a special edge in $G^{\prime}$.
@@
For each $u\neq v\in V(G)$, it follows from the inductive hypothesis that $\mathcal{H}_{u}$ induces a connected subgraph of $Q^{*}$. It remains to show that $\mathcal{H}_{v}$ induces a connected subgraph of $Q^{*}$. Let $C$ denote the cycle $(u_{0},\ldots,u_{\deg(v)-1},u_{0})$ added in VB$(v)$. Since we assumed (by Proposition 27) that $\mathcal{H}$ has no containments, it follows that $\cup_{i=0}^{\deg(v)-1}\mathcal{H}^{\prime}_{u_{i}}=\mathcal{H}_{v}$ as there is no subgraph containing only the vertex $v$. If none of the edges of $C$ are in $\textrm{Spl}_{\mathcal{H}^{\prime}}(E)$, then $\mathcal{H}_{v}$ is connected since adjacent vertices of $C$ share at least one subgraph and by our assumption on applying $\textsc{VB}(v)$, $\mathcal{H}^{\prime}_{u_{i}}\neq\emptyset$ for any $i=\{0,\ldots,deg(v)-1\}$. On the other hand, if an edge $e=\{u_{i},u_{i+1}\}$ (where indices are taken $\mathrm{mod}\deg(v)$) of $C$ is in $\textrm{Spl}_{\mathcal{H}^{\prime}}(E)$, by the inductive hypothesis, since $Q^{*}$ satisfies the special edge property, at least one subgraph from $\mathcal{H}^{\prime}_{u_{i}}$ and one subgraph from $\mathcal{H}^{\prime}_{u_{i+1}}$ are adjacent in $Q^{*}$. Since $\cup_{i=0}^{\deg(v)-1}\mathcal{H}^{\prime}_{u_{i}}=\mathcal{H}_{v}$, and $C$ is a cycle, it follows that $\mathcal{H}_{v}$ is connected and thus $Q^{*}$, is the desired dual support for $(G,\mathcal{H})$ that also satisfies the special edge property. ∎
@@ 6.3 Intersection Support
Now, we apply vertex bypassing to obtain a graph where none of the $\mathcal{K}$-vertices are maximal. For a $\mathcal{K}$-vertex $v$, if an edge $e\sim v$ is such that $\mathcal{K}_{v}=\mathcal{K}_{e}$, we say that $e$ is full for $v$. If a $\mathcal{K}$-vertex does not have a full edge incident on it, then we say that it is maximal w.r.t. $\mathcal{K}$ subgraphs. In this case, $\mathcal{K}_{e}\subset\mathcal{K}_{v}$ for all ${e}\sim v$. In the following, we repeatedly apply vertex bypassing to a maximal $\mathcal{K}$-vertex of maximum depth until no $\mathcal{K}$-vertex is maximal. Note that a maximum depth $\mathcal{K}$-vertex need not be maximal.
@@
Let $(G,\mathcal{H},\mathcal{K})$ be a cross-free intersection system of genus $g$. Then, we can modify the arrangement to a cross-free arrangement $(G^{\prime},\mathcal{H},\mathcal{K}^{\prime})$ so that $G^{\prime}$ has genus $g$, no $\mathcal{K}$-vertex of $G^{\prime}$ is maximal, and an intersection support for $(G^{\prime},\mathcal{H},\mathcal{K}^{\prime})$ is also an intersection support for $(G,\mathcal{H},\mathcal{K})$.
@@
We assume that a cross-free embedding of $(G,\mathcal{H},\mathcal{K})$ is given, and with a slight abuse of notation, we use $G$ to also refer to the embedded graph. Let $d$ denote the maximum depth of a maximal $\mathcal{K}$-vertex, and let $n_{d}$ denote the number of maximal $\mathcal{K}$-vertices of depth $d$. We repeatedly choose a maximal $\mathcal{K}$-vertex $v$ of depth $d$ and apply VB$(v)$. The operation VB$(v)$ modifies the graph $G$ and the subgraphs in $\mathcal{K}$, but does not modify any subgraph in $\mathcal{H}$ as $v$ is a $\mathcal{K}$-vertex. Let $v$ be a $\mathcal{K}$-vertex to which we apply $\textsc{VB}(v)$. Let $u_{0},\ldots,u_{deg(v)-1}$ be the new vertices added corresponding to the edges $\{v,v_{i}\}$ in $G$ for $10\leq i\leq deg(v)-$ 1. The vertices $u_{i}$, $i=0,\ldots,deg(v)-1$ are also $\mathcal{K}$-vertices. Since the edge $\{u_{i},v_{i}\}$ is full for $u_{i}$, the vertex $u_{i}$ is not maximal w.r.t. $\mathcal{K}$. Hence, none of the newly added vertices are maximal $\mathcal{K}$-vertices.
Let $(G^{\prime},\mathcal{H},\mathcal{K}^{\prime})$ denote the resulting intersection system. Let $d^{\prime}$ and $n_{d^{\prime}}$ denote respectively, the maximum depth of a maximal $\mathcal{K}$-vertex, and the number of maximal $\mathcal{K}$-vertices in $(G^{\prime},\mathcal{H},\mathcal{K}^{\prime})$ of depth $d^{\prime}$. It follows that $(d^{\prime},n_{d^{\prime}})$ is lexicographically smaller than $(d,n_{d})$ and hence the process eventually stops. Since the newly added vertices and edges are not contained in any subgraph $H\in\mathcal{H}$, it follows that $\mathcal{H}$ remains cross-free in $G^{\prime}$. Further, the fact that subgraphs in $\mathcal{K}^{\prime}$ remain connected and cross-free follows from Lemmas 22 and 23. The fact that $G^{\prime}$ has genus $g$ follows from the fact that the operation of vertex bypassing preserves the embedding of the resulting graph on the same surface as that of the original graph. Thus, $(G^{\prime},\mathcal{H},\mathcal{K}^{\prime})$ is cross-free. Since the underlying hypergraph (,$(\mathcal{H},\{\mathcal{H}_{K}\}_{K\in\mathcal{K}})$ is not modified, an intersection support for $(G^{\prime},\mathcal{H},\mathcal{K}^{\prime})$ is also an intersection support for $(G,\mathcal{H},\mathcal{K})$. ∎
@@
By Lemma 30, it is sufficient to prove an intersection support for $(G,\mathcal{H},\mathcal{K})$ when no $\mathcal{K}$-vertex is maximal. At each $\mathcal{K}$-vertex $v$, we add a dummy subgraph $F_{v}$. Let $\mathcal{F}$ denote the set of dummy subgraphs thus added. For the intersection system $(G,\mathcal{H}\cup\mathcal{F},\mathcal{K})$, we can obtain an intersection support $Q$ by applying Lemma 29. We now show that we can obtain an intersection support for $(G,\mathcal{H},\mathcal{K})$ from $Q$.
@@
If $G$ does not contain any $\mathcal{K}$-vertex, then by Lemma 29, we are done. So, suppose that there are some $\mathcal{K}$ vertices in $G$. By Lemma 30, we can assume without loss of generality that no $\mathcal{K}$-vertex is maximal.
At each $\mathcal{K}$-vertex $v$ of $G$, we add a dummy subgraph $F_{v}$. Let $\mathcal{F}$ denote the set of dummy subgraphs added. Consider the intersection system $(G,\mathcal{H}\cup\mathcal{F},\mathcal{K})$. By Lemma 29, there is a dual support $Q$ for $(G,\mathcal{H}\cup\mathcal{F})$ with special edge property such that $Q$ is an intersection support for $(G,\mathcal{H}\cup\mathcal{F},\mathcal{K})$ and $Q$ has genus at most $g$. Since $Q$ is an intersection support, each $K\in\mathcal{K}$ induces a connected subgraph $\mathcal{H}_{K}$ of $Q$. That is, in the graph system $(Q,\mathcal{K})$, each $K\in\mathcal{K}$ induces a connected subgraph of $Q$. Abusing notation, we use the names of the subgraphs in $\mathcal{H}\cup\mathcal{F}$ to denote their corresponding vertices in $Q$. Consider a dummy subgraph $F_{u}$ corresponding to a $\mathcal{K}$-vertex $u$. Since $u$ is not maximal in $G$ w.r.t. $\mathcal{K}$, it has a full edge $e=\{u,v\}$ incident to it, i.e., $\mathcal{K}_{u}=\mathcal{K}_{e}$. But, this implies $\mathcal{K}_{u}\subseteq\mathcal{K}_{v}$. Note that all edges incident to $u$ in $G$ are present in the dual support $Q$ as special edges since the dummy subgraph $F_{u}$ consists of a single vertex $u$. Thus, the edge $e$ is incident to $F_{u}$ in $Q$. If $X\in\mathcal{H}\cup\mathcal{F}$ is the other end of $e$ in $Q$, then $\mathcal{K}_{F_{u}}\subseteq\mathcal{K}_{X}$ in the graph system $(Q,\mathcal{K})$ i.e., $F_{u}$ is not a maximal vertex in $(Q,\mathcal{K})$. Hence, none of the dummy subgraphs $\mathcal{F}$ corresponds to a maximal vertex in $(Q,\mathcal{K})$.
Now we assign a 2-coloring to the vertices of $Q$. Let $c:V(Q)\to\{\mathbf{r},\mathbf{b}\}$ be such that $c(F)=\mathbf{r}$ for each $F\in\mathcal{F}$, and $c(H)=\mathbf{b}$ for each $H\in\mathcal{H}$. Now, $Q$ satisfies the conditions of Lemma 26 that no red vertex is maximal and each $K\in\mathcal{K}$ induces a connected subgraph of $Q$. Therefore, the graph system $(Q,\mathcal{K})$ admits a support $\tilde{Q}$ on $\mathcal{H}$ of genus at most $g$. Since each $K\in\mathcal{K}$ induces a connected subgraph in $\tilde{Q}$, we have the desired support. ∎

### v3

@@ 6.2 Dual Support
In this section, we show that for a cross-free system on a graph of genus $g$, there is a dual support of genus at most $g$. First, we show that for constructing a dual support, we can assume that there are no two $H,H^{\prime}\in\mathcal{H}$ such that $V(H)\subseteq V(H^{\prime})$.
@@
Let $(G,\mathcal{H})$ be a graph system. Let $\mathcal{H}^{\prime}\subseteq\mathcal{H}$ be maximal such that $\forall\;H,H^{\prime}\in\mathcal{H}^{\prime}$, $H\setminus H^{\prime}\neq\emptyset$, and that for each $\in\mathcal{H}\setminus\mathcal{H}^{\prime}, HH$$\subseteq H^{\prime}$ for some $H^{\prime}\in\mathcal{H}^{\prime}$. If $Q^{\prime}$ is a dual support for $(G,\mathcal{H}^{\prime})$ then, there is a dual support $Q^{*}$ for $(G,\mathcal{H})$ such that if $Q^{\prime}$ has genus $g$, then $Q^{*}$ has genus $g$.
@@
Consider the containment order $\mathcal{P}=(\mathcal{H},\preceq)$ where for $^{\prime}\in\mathcal{H}, HHH\preceq H^{\prime}$ iff $V(H)\subseteq V(H^{\prime})$. Then $\mathcal{P}$ is a partial ordered set. Let $\ell(\mathcal{P})$ be the maximum length of a chain in $\mathcal{P}$, and $c(\mathcal{P})$ be the number of chains of length $\ell(\mathcal{P})$. We prove by induction on $(\ell(\mathcal{P}),c(\mathcal{P}))$, that a dual support $Q^{\prime}$ for a lexicographically smaller system can be extended to a dual support $Q^{*}$ of a larger system so that $Q^{*}$ has the same genus as $Q^{\prime}$. If we show this, then we can remove all containments, construct a support, and extend it to obtain a dual support for the original system without increasing the genus.
@@
Consider a longest chain $C$ in $\mathcal{P}$, and let $H$ be the minimum element in $C$. Let $\mathcal{H}^{\prime}=\mathcal{H}\setminus\{H\}$. Then, for containment order $\mathcal{P}^{\prime}$ on $\mathcal{H}^{\prime}$, $(\ell(\mathcal{P}^{\prime}),c(\mathcal{P}^{\prime}))$ is lexicographically smaller than $(\ell(\mathcal{P}),c(\mathcal{P}))$. Therefore, by the inductive hypothesis, the statement of the proposition holds for $(G,\mathcal{H}^{\prime})$, and let $Q^{\prime}$ denote its dual support. To obtain a dual support $Q^{*}$ for $(G,\mathcal{H})$, we add a vertex corresponding to $H$. Let $H^{\prime}$ be an immediate successor of $H$ in $\mathcal{P}$ arbitrarily chosen. Connect $H$ to $H^{\prime}$. Since we added a new vertex of degree 1, it follows that if $Q^{\prime}$ has genus $g$, then $Q^{*}$ has genus $g$.
@@
For each vertex $v\in V(G)$, the subgraphs containing $v$ trivially induce a connected subgraph in $Q^{*}$ since $0pt(v)=1$ for all $v\in V(G)$. Consider a special edge $e=\{u,v\}\in E(G)$. Since both end-points belong to a subgraph in $\mathcal{H}$, and $\mathcal{H}_{u}\cap\mathcal{H}_{v}=\emptyset$, it implies that $\mathcal{H}_{u}\not\subseteq\mathcal{H}_{v}$ and $\mathcal{H}_{v}\not\subseteq\mathcal{H}_{u}$. In the construction of $Q^{*}$ therefore, $e$ is not contracted. Further, $u$ and $v$ are contracted to vertices corresponding to the unique subgraph containing $u$ and $v$, respectively. Therefore, $Q^{*}$ satisfies the special edge property. ∎
@@
Therefore, we can assume that for a maximum depth vertex $v$ and each $e\sim v$, $\mathcal{H}_{e}\subset\mathcal{H}_{v}$. We apply VB(v)(v) to obtain the system $(G^{\prime},\mathcal{H}^{\prime})$. By Lemma  25, the new vertices $u_{0},\ldots,u_{deg(v)-1}$ in $G^{\prime}$ obtained on applying Vertex Bypassing to $v$ have depth at most d1d-1. Hence, $(d^{\prime},n_{d^{\prime}})\prec(d,n_{d})$, where $d^{\prime}$ and $n_{d^{\prime}}$ are respectively, the depth of a maximum depth vertex, and the number of vertices of maximum depth in $(G^{\prime},\mathcal{H}^{\prime})$. We can assume that $\mathcal{H}^{\prime}_{u_{i}}\neq\emptyset$ for each $i\in\{0,\ldots,\deg(v)-1\}$. Otherwise, we can contract the edge $\{u_{i},u_{i+1}\}$ (with indices taken $\mathrm{mod}$ $\deg(v)$) as this does not violate the cross-free condition on $\mathcal{H}^{\prime}$. Further, there is an injective correspondence between the special edges in $G$ and the special edges in $G^{\prime}$: For $i=0,\ldots,deg(v)-1$, if $\{v,v_{i}\}$ in $G$ is a special edge, the corresponding edge $\{u_{i},v_{i}\}$ is a special edge in $G^{\prime}$.
@@
For each $u\neq v\in V(G)$, it follows from the inductive hypothesis that $\mathcal{H}_{u}$ induces a connected subgraph of $Q^{*}$. It remains to show that $\mathcal{H}_{v}$ induces a connected subgraph of $Q^{*}$. Let $C$ denote the cycle $(u_{0},\ldots,u_{\deg(v)-1},u_{0})$ added in VB(v)(v). Since we assumed (by Proposition 27) that $\mathcal{H}$ has no containments, it follows that $\cup_{i=0}^{\deg(v)-1}\mathcal{H}^{\prime}_{u_{i}}=\mathcal{H}_{v}$ as there is no subgraph containing only the vertex $v$. If none of the edges of $C$ are in $\textrm{Spl}_{\mathcal{H}^{\prime}}(E)$, then $\mathcal{H}_{v}$ is connected since adjacent vertices of $C$ share at least one subgraph and by our assumption on applying $\textsc{VB}(v)$, $\mathcal{H}^{\prime}_{u_{i}}\neq\emptyset$ for any $i=\{0,\ldots,deg(v)-1\}$. On the other hand, if an edge $e=\{u_{i},u_{i+1}\}$ (where indices are taken $\mathrm{mod}\deg(v)$) of $C$ is in $\textrm{Spl}_{\mathcal{H}^{\prime}}(E)$, by the inductive hypothesis, since $Q^{*}$ satisfies the special edge property, at least one subgraph from $\mathcal{H}^{\prime}_{u_{i}}$ and one subgraph from $\mathcal{H}^{\prime}_{u_{i+1}}$ are adjacent in $Q^{*}$. Since $\cup_{i=0}^{\deg(v)-1}\mathcal{H}^{\prime}_{u_{i}}=\mathcal{H}_{v}$, and $C$ is a cycle, it follows that $\mathcal{H}_{v}$ is connected and thus $Q^{*}$, is the desired dual support for $(G,\mathcal{H})$ that also satisfies the special edge property. ∎
By the special edge property of a dual support, we make use of the following observation for the construction of an intersection support in the following section.
Observation 29.
Given a graph system $(G,\mathcal{H})$, let $Q^{*}$ be a dual support with the special edge property. Let $H_{u}=\{u\}$ and $H_{v}=\{v\}$ be subgraphs in $\mathcal{H}$ for some depth 1 vertices $u,v\in V(G)$. Then we can assume wlog that $u$ and $v$ are adjacent in $G$ if and only if $H_{u}$ and $H_{v}$ are adjacent in $Q^{*}$. Indeed, if $u,v$ are adjacent in $G$, then $H_{u},H_{v}$ should be adjacent in $Q^{*}$ by the special edge property. If $u,v$ are not adjacent in $G$ but $H_{u},H_{v}$ are adjacent in $Q^{*}$, then removing the edge $\{H_{u},H_{v}\}$ from $Q^{*}$ does not violate the conditions of a dual support with the special edge property.
@@ 6.3 Intersection Support
Now, we apply vertex bypassing to obtain a graph system where none of the $\mathcal{K}$-vertices are maximal. For a $\mathcal{K}$-vertex $v$, if an edge $e\sim v$ is such that $\mathcal{K}_{v}=\mathcal{K}_{e}$, we say that $e$ is full for $v$. If a $\mathcal{K}$-vertex does not have a full edge incident on it, then we say that it is maximal w.r.t. $\mathcal{K}$ subgraphs. In this case, $\mathcal{K}_{e}\subset\mathcal{K}_{v}$ for all ${e}\sim v$. Note that a maximum depth $\mathcal{K}$-vertex need not be maximal. In the following, we repeatedly apply vertex bypassing to a maximal $\mathcal{K}$-vertex of maximum depth until no $\mathcal{K}$-vertex is maximal. However, we require an additional property like in the construction of the primal support, that there are no adjacent twins. We recall the definition of twins and adjacent twins here. If two $\mathcal{K}$-vertices $u,v$ of $G$ are contained in the same set of subgraphs of $\mathcal{K}$, they are said to be twins. If, in addition, $u$ and $v$ are adjacent, then they are said to be adjacent twins.
The proof of the lemma below follows from arguments similar to those of Theorem 7.
@@
Let $(G,\mathcal{H},\mathcal{K})$ be a cross-free intersection system of genus $g$. Then, we can modify the arrangement to a cross-free arrangement $(G^{\prime},\mathcal{H},\mathcal{K}^{\prime})$ such that (i) $G^{\prime}$ has genus $g$, (ii) no $\mathcal{K}$-vertex in $(G^{\prime},\mathcal{H},\mathcal{K}^{\prime})$ is maximal, (iii) no two $\mathcal{K}$-vertices are adjacent twins, and (iv) an intersection support for $(G^{\prime},\mathcal{H},\mathcal{K}^{\prime})$ is also an intersection support for $(G,\mathcal{H},\mathcal{K})$.
@@
We assume that a cross-free embedding of $(G,\mathcal{H},\mathcal{K})$ is given, and with a slight abuse of notation, we use $G$ to also refer to the embedded graph. We will modify the cross-free graph system $(G,\mathcal{H},\mathcal{K})$ to a cross-free graph system $(G^{\prime},\mathcal{H},\mathcal{K}^{\prime})$ such that $G^{\prime}$ has an embedding on the same surface as $G$ does, and that the intersection hypergraph defined by $(G,\mathcal{H},\mathcal{K})$ is isomorphic to that defined by $(G^{\prime},\mathcal{H},\mathcal{K}^{\prime})$. This will prove $(i)$ and $(iv)$. Also, in the process of obtaining $(G^{\prime},\mathcal{H},\mathcal{K}^{\prime})$, as long as there are adjacent twins among $\mathcal{K}$-vertices, we contract them to a single vertex. Thus, we will be done with condition $(iii)$ as well. So, in the rest of the proof, we claim to establish condition $(ii)$ of the lemma.
Let $d$ denote the maximum depth of a maximal $\mathcal{K}$-vertex in $G$. We prove the argument by induction on $d$. For $i=1,2,\ldots,d$, let $S_{i}$ denote the set of maximal $\mathcal{K}$-vertices of depth $i$. Suppose $d=1$, and consider any $v\in S_{1}$. By the definition of a maximal vertex w.r.t. $\mathcal{K}$, the subgraph $K\in\mathcal{K}$ containing $v$, does not contain any other vertex of $G$. Thus $\mathcal{H}_{K}=\emptyset$. Let $\mathcal{K}_{1}\subseteq\mathcal{K}$ denote the collection of all subgraphs consisting of single vertices in $S_{1}$, and let $\mathcal{K}^{\prime}=\mathcal{K}\setminus\mathcal{K}_{1}$. Then $(G^{\prime},\mathcal{H},\mathcal{K}^{\prime})$ is the required graph system, where $G^{\prime}=G$.
Let our claim be true for $d<k$ for some positive integer $k\geq 2$. Consider a cross-free graph system $(G,\mathcal{H},\mathcal{K})$ and a 2-coloring $c:V(G)\to\{\mathbf{r},\mathbf{b}\}$ such that $d=k$. We apply vertex bypassing to all the vertices in $S_{k}$. Note that this does not modify any subgraph in $\mathcal{H}$. For any pair of adjacent twins created among the $\mathcal{K}$-vertices, we contract them to a single $\mathcal{K}$-vertex until there is no such twin. Let $(G^{\prime\prime},\mathcal{H},\mathcal{K}^{\prime\prime})$ denote the resulting graph system. By Lemma 23 and the contraction among the adjacent $\mathcal{K}$-twins, it follows that $(G^{\prime\prime},\mathcal{K}^{\prime\prime})$ remains cross-free. Since for any vertex $v\in V(G)$ that lies in a subgraph $H\in\mathcal{H}$, the cyclic order of its neighbours is preserved in $G^{\prime\prime}$, the graph system $(G^{\prime\prime},\mathcal{H})$ is cross-free. Therefore, $(G^{\prime\prime},\mathcal{H},\mathcal{K}^{\prime\prime})$ is a cross-free intersection system. By the construction of $G^{\prime\prime}$, note that a subgraph $H\in\mathcal{H}$ intersects a subgraph $K\in\mathcal{K}$ in $G$ if and only if $H$ intersects in $G^{\prime\prime}$ with the representative $K^{\prime\prime}$ of $K$. It follows that the intersection hypergraphs defined by the two intersection systems are isomorphic.
By an argument similar to that in the proof of Theorem 7, the maximum depth of a maximal $\mathcal{K}$-vertex in $(G^{\prime\prime},\mathcal{H},\mathcal{K}^{\prime\prime})$ is at most k1k-1. By the induction hypothesis, $(G^{\prime\prime},\mathcal{H},\mathcal{K}^{\prime\prime})$ can be modified to a graph system $(G^{\prime},\mathcal{H},\mathcal{K}^{\prime})$ satisfying properties (i)(i)-$(iv)$ of the lemma. $(G^{\prime},\mathcal{H},\mathcal{K}^{\prime})$ is the required cross-free intersection system since the intersection hypergraph defined by $(G^{\prime},\mathcal{H},\mathcal{K}^{\prime})$ is isomorphic to that defined by $(G^{\prime\prime},\mathcal{H},\mathcal{K}^{\prime\prime})$, which itself is isomorphic to the intersection hypergraph defined by $(G,\mathcal{H},\mathcal{K})$. ∎
@@
By Lemma 31, it is sufficient to prove an intersection support for $(G,\mathcal{H},\mathcal{K})$ when no $\mathcal{K}$-vertex is maximal, and there is no adjacent twin among $\mathcal{K}$-vertices. At each $\mathcal{K}$-vertex $v$, we add a dummy subgraph $F_{v}$. Let $\mathcal{F}$ denote the set of dummy subgraphs thus added. For the intersection system $(G,\mathcal{H}\cup\mathcal{F},\mathcal{K})$, we can obtain an intersection support $Q$ by applying Lemma 30. We now show that we can obtain an intersection support for $(G,\mathcal{H},\mathcal{K})$ from $Q$ using Lemma 26 and Observation 29.
@@
If $G$ does not contain any $\mathcal{K}$-vertex, then by Lemma 30, we are done. So, suppose that there are some $\mathcal{K}$-vertices in $G$. By Lemma 31, we can assume without loss of generality that no $\mathcal{K}$-vertex is maximal, and that there are no adjacent twins among the $\mathcal{K}$-vertices.
At each $\mathcal{K}$-vertex $v$ of $G$, we add a dummy subgraph $F_{v}$. Let $\mathcal{F}$ denote the set of dummy subgraphs added. Let $\mathcal{H}^{\prime}=\mathcal{H}\cup\mathcal{F}$ and consider the intersection system $(G,\mathcal{H}^{\prime},\mathcal{K})$. By Lemma 30, there is a dual support $Q$ for $(G,\mathcal{H}^{\prime})$ with the special edge property such that $Q$ is an intersection support for $(G,\mathcal{H}^{\prime},\mathcal{K})$, and $Q$ has genus at most $g$. Since $Q$ is an intersection support, each $K\in\mathcal{K}$ induces a connected subgraph $\mathcal{H}^{\prime}_{K}$ of $Q$. Abusing notation, we use the labels of the subgraphs in $\mathcal{H}\cup\mathcal{F}$ to denote their corresponding vertices in $Q$. Consider a dummy subgraph $F_{u}$ corresponding to a $\mathcal{K}$-vertex $u$. Since $u$ is not maximal in $G$ w.r.t. $\mathcal{K}$, it has a full edge $e=\{u,v\}$ incident to it, i.e., $\mathcal{K}_{u}=\mathcal{K}_{e}$. But, this implies $\mathcal{K}_{u}\subseteq\mathcal{K}_{v}$. Note that for every edge incident to $u$ in $G$, there is a corresponding edge in the dual support $Q$ as a special edge incident to the dummy subgraph $F_{u}$, since $F_{u}$ consists of a single vertex $u$, and $|\mathcal{H}^{\prime}_{u}|=1$. In particular, let $e^{\prime}$ be the copy of $e$ incident to $F_{u}$ in $Q$. If $X\in\mathcal{H}\cup\mathcal{F}$ is the other end of $e^{\prime}$, then $\mathcal{K}_{F_{u}}\subseteq\mathcal{K}_{X}$ in the graph system $(Q,\mathcal{K})$ i.e., $F_{u}$ is not a maximal vertex in $(Q,\mathcal{K})$. Hence, none of the dummy subgraphs $\mathcal{F}$ corresponds to a maximal vertex in $(Q,\mathcal{K})$.
Now we assign a 2-coloring to the vertices of $Q$. Let $c:V(Q)\to\{\mathbf{r},\mathbf{b}\}$ be such that $c(F)=\mathbf{r}$ for each $F\in\mathcal{F}$, and $c(H)=\mathbf{b}$ for each $H\in\mathcal{H}$. To complete the argument, we need to show that there are no adjacent red twins in the graph system $(Q,\mathcal{K})$. Consider any two red vertices $F_{u}$ and $F_{v}$ of $Q$, corresponding to some pair of vertices $u$ and $v$ of G, respectively. Since $|\mathcal{H}_{u}|=1=|\mathcal{H}_{v}|$, by Observation 29, we can assume that $F_{u}$ and $F_{v}$ are adjacent in $Q$ if and only if $u$ and $v$ are adjacent in $G$. If $F_{u}$ and $F_{v}$ are not adjacent, then trivially they are not adjacent red twins. So, suppose they are adjacent. Then $u$ and $v$ are also adjacent in $G$. By assumption, $G$ does not contain adjacent $\mathcal{K}$-vertices that are twins, thus $\mathcal{K}_{u}\neq\mathcal{K}_{v}$. It follows that in the graph system $(Q,\mathcal{K})$, we have $\mathcal{K}_{F_{u}}\neq\mathcal{K}_{F_{v}}$, and hence, $F_{u}$ and $F_{v}$ are not twins in $Q$. Therefore, $Q$ satisfies the conditions of Lemma 26 that no red vertex is maximal and each $K\in\mathcal{K}$ induces a connected subgraph of $Q$. Therefore, the graph system $(Q,\mathcal{K})$ admits a support $\tilde{Q}$ on $\mathcal{H}$ of genus at most $g$. Since each $K\in\mathcal{K}$ induces a connected subgraph in $\tilde{Q}$, we have the desired intersection support for $(G,\mathcal{H},\mathcal{K})$. ∎